### 12 Jan 2024 v1

# Heterogeneous Low-Rank Approximation for Federated Fine-tuning of On-Device Foundation Models

Abstract
Large foundation models (FMs) adapt surprisingly well to specific domains or tasks with fine-tuning. Federated learning (FL) further enables private FM fine-tuning using the local data on devices. However, the standard FMs’ large size poses challenges for resource-constrained and heterogeneous devices. To address this, we consider FMs with reduced parameter sizes, referred to as on-device FMs (ODFMs). While ODFMs allow on-device inference, computational constraints still hinder efficient federated fine-tuning. We propose a parameter-efficient federated fine-tuning method for ODFMs using heterogeneous low-rank approximations (LoRAs) that addresses system and data heterogeneity. We show that homogeneous LoRA ranks face a trade-off between overfitting and slow convergence, and propose HetLoRA, which employs heterogeneous ranks across clients and eliminates the shortcomings of homogeneous LoRA. By applying rank self-pruning locally and sparsity-weighted aggregation at the server, we combine the advantages of high and low-rank LoRAs, which achieves improved convergence speed and final performance compared to homogeneous LoRA. Furthermore, it offers enhanced computation efficiency compared to full fine-tuning, making it suitable for heterogeneous devices while preserving data privacy.
Heterogeneous Low-Rank Approximation for Federated Fine-tuning of On-Device Foundation Models
[…]
1 Introduction
The emerging foundation models (FMs) Bommasani et al. (2022); Zhou et al. (2023); Radford et al. (2021); Devlin et al. (2019); OpenAI (2023); Google (2022); Touvron et al. (2023); Brown et al. (2020); Google (2022); Driess et al. (2023); Google (2023) have shown remarkable zero/few shot learning capabilities, performing well on a variety of tasks including text/image generation with prompts, language translation, solving math problems, and conversing in natural language. Standard FMs, however, demand costly resources for directly fine-tuning their entire parameter space. To tackle this issue, many recent works have proposed different parameter-efficient fine-tuning (PEFT) methods of FMs such as prompt tuning Lester et al. (2021), utilizing adapters Houlsby et al. (2019), or low-rank adaptation (LoRA) of the original model Hu et al. (2021) which freezes the original pre-trained parameters of the FM and train additional, smaller number of parameters instead.
These PEFT methods, however, assume that i) FMs are deployed to and trained with the data of a single machine/client for adaptation to the downstream task and that ii) the client has enough resources to even fit a standard FM of hundred billion size for, at least, inference. In practice, there are frequently cases where we are interested in fine-tuning FMs for on-device private data that is distributed across multiple devices (clients). For instance, sensitive and private data such as medical information or law-related documents may be hard to collect centrally in a private manner and fine-tuning of the FMs may need to be done at the edge Manoel et al. (2023); Shoham and Rappoport (2023); Zhang et al. (2023c).
In our work, we focus on federated fine-tuning, which we define as training a set of parameters collaboratively across clients to obtain a global set of parameters that can be plugged in to the FM for the targeted downstream task. Note that federated fine-tuning is orthogonal to personalization of FMs in federated learning (FL) Guo et al. (2023), which aims to train parameters that perform well for individual clients rather than general downstream tasks. We also define on-device FMs (ODFMs) as models with fewer than several billion parameters that are able to fit into memory on limited capacity clients considering current hardwares.
Zero-Shot Few-Shot Full-Training
PaLM2 XXS $2930.23$ $2541.86$ $\mathbf{23.71}$
[…]
Table 1: Perplexity of PaLM2 for zero-shot, few-shot (5 communication rounds), and full federated fine-tuning (200 communication rounds) for chat response on the multi-session chat data (further experimental details are in Section 4.)
$r=1$ $r=5$ $r=10$ $r=20$ $r=50$ $r=100$ $r=150$ $r=200$
PaLM2 XXS, PaLM2 XS $0.02\%$ $0.11\%$ $0.21\%$ $0.42\%$ $1.05\%$ $2.10\%$ $3.14\%$ $4.19\%$
Table 2: Percentage of the LoRA parameters’ size for different ranks $r$ compared to the original pre-trained ODFM’s parameter size. Even for large ranks such as $r=200$ the trainable LoRA parameters’ size compared to the original pre-trained ODFM size is less than $5\%$ for both PaLM2-XS and PaLM2-XXS.
Federated fine-tuning of ODFMs entails unique challenges that are not present in either standard PEFT of FMs or the standard federated training of models that are not FMs. First, FMs have their zero/few-shot learning capability often supported by their large parameter space that is trained on massive data. However, as shown in Table 1 and previous literature Kojima et al. (2022), FMs’ performance can deteriorate as their sizes get smaller and federated fine-tuning may not merely be useful but even inevitable for ODFMs to perform well for downstream tasks on devices.
Moreover, devices have limited and heterogeneous system capabilities Wang et al. (2019); Bonawitz et al. (2016) and data-heterogeneity Sahu et al. (2020). A suitable PEFT method that flexibly adapts to the heterogeneous devices should be devised for federated fine-tuning of ODFMs. Previous work evaluated PEFT with FL via performing a general evaluation over different PEFT methods naïvely combined with FL Guo et al. (2022); Zhang et al. (2023d); Chen et al. (2022); Wortsman et al. (2023); Yu et al. (2023). However, they do not investigate the practical setting for ODFMs where PEFT methods are catered to heterogeneous system-capabilities and data such as by applying different PEFT strategies to different clients.
In our work. we focus on one of the most prominent PEFT methods, LoRA which proposes to train low-rank approximations of the original model. Using LoRA, the number of trainable parameters is greatly reduced to at most $0.02\%$ of the original ODFM size (see Table 2). The simplest way to apply LoRA to federated fine-tuning is training with homogeneous rank $r$ across the clients as one would train any global model with FL. However, this does not cater to the heterogeneity in FL, where it is even difficult to choose the right LoRA rank for resource limited mobile devices with natural system and data heterogeneity.
To this end, we propose heterogeneous LoRA, namely HetLoRA in short, for federated fine-tuning to cater to system and data heterogeneity and outperform the naïve combination of LoRA and federated fine-tuning where homoegneous ranks are applied across clients. We show the performance of PaLM 2 Google (2023) of XXS and XS size for chat responses on the multi-session chat data Xu et al. (2021) and text summarization for the Reddit data Völske et al. (2017), both which are real world data from clients. Our contributions can be summarized as follows:
  • We propose HetLoRA that can apply different rank LoRA modules to different clients to cater to the heterogeneous system capabilities and data complexities of the clients, via utilizing rank self-pruning and sparsity-weighted aggregation.
  • We show the performance of naïvely applying LoRA with homogeneous ranks across clients for federated fine-tuning, and show that while large ranks help in speeding-up training, they lead to faster overfitting while smaller ranks are slower in training but does not suffer from overfitting.
  • We then evaluate HetLoRA to show that it outperforms naïvely applying homogeneous ranks across clients in terms of both training speed, communication/computation efficiency, and final performance, gaining the best of both worlds of homogeneous LoRA with high and low ranks.
2 Related Work
Parameter-Efficient Fine Tuning. There has been a plethora of recent work on PEFT which either trains a subset of parameters within the existing FM whilst other parameters are freezed or introduces an additional set of trainable parameters whilst keeping the original FM freezed. For the former, methods such as head or bias fine-tuning Wei et al. (2021); Bu et al. (2022); Lee et al. (2019); Zaken et al. (2021) has been explored, while for the latter, methods such as adapters Houlsby et al. (2019), prompt Lester et al. (2021) or prefix-tuning Li and Liang (2021), and low-rank approximation Hu et al. (2021) has been proposed. While these number of methods has been proven to perform as well as full model fine-tuning with just a few number of parameters for the centralized setting, it has not been thoroughly explored how these methods perform for a much smaller FM such as ODFMs, in the decentralized setting where clients’ system-capacities can be heterogeneous and much limited.
Federated Fine-Tuning. Recently, interest in the intersection of foundation models and FL has notably increased Zhou et al. (2023); Yu et al. (2023). Many recent work has proposed to simply combine the PEFT methods devised for the centralized setting to FL such as training prompts or adapters collaboratively with FL Guo et al. (2022); Chen et al. (2022); Zhang et al. (2023a); Shysheya et al. (2023); Legate et al. (2023). Another line of work has proposed to perform a few-shot or nearly zero-shot training of FMs with FL for improved communication-efficiency Wortsman et al. (2023); Zhang et al. (2023d). However, these work either overlooks that most devices do not have the resource to fit a general FM Touvron et al. (2023); Brown et al. (2020) (>8B parameters) even for inference or does not consider the heterogeneous system capacities of the clients. It is detrimental to consider these factors since FMs that can actually fit to the devices in FL are usually much smaller making them weaker in the general intelligence capabilities, and also heterogeneous system capacities may prohibit deploying the same sized PEFT parameters across clients.
Only a few number of recent work has looked in to using LoRA for FL. For instance, in Babakniya et al. (2023), the importance of the initialization for the LoRA modules is evaluated where they propose to train the LoRA modules with FL and then perform singular value decomposition (SVD) to gain a good initialization of the LoRA modules. However, the training process of LoRA itself is not altered to adapt to heterogeneous system capabilieis of devices. Another recent work Yi et al. (2023) has evaluated LoRA in the context of personalized FL, but other than applying LoRA to personalization, the LoRA method itself is, again, not changed. Our work proposes heterogeneous LoRA for federated fine-tuning where heterogeneous ranks are deployed and trained across clients by a new algorithm that includes rank self-pruning and sparsity weighted aggregation.
3 Federated Fine-Tuning with LoRA
3.1 Preliminaries
Formally, we define the pre-trained ODFM as $\mathbf{W}_{0}\in\mathbb{R}^{d\times l}$ and the trainable low-rank decomposed matrix as $\Delta\mathbf{W}\in\mathbb{R}^{d\times l}$. In standard LoRA Hu et al. (2021), the low-rank decomposition of $\Delta\mathbf{W}$ is constructed such that $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times l}$ are the low rank decomposition of $\Delta\mathbf{W}$ with identical rank $r$ in the centralized setting. Now, let us consider LoRA for federated fine-tuning where there are $M$ total clients. Each client $k\in[M]$ has private data $\mathcal{B}_{k}$ and its corresponding local empirical loss function $F_{k}(\mathbf{W})=\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell(\mathbf{W},\xi)$, where $\ell(\mathbf{W},\xi)$ is the loss for model $\mathbf{W}$ at data sample $\xi$. The optimization task for federated fine-tuning is to collaboratively find the global parameters which we define as $\overline{\mathbf{B}}$ and $\overline{\mathbf{A}}$, given the pretrained knowledge $\mathbf{W}_{0}$ that can minimize the global objective $F(\overline{\mathbf{W}})=\frac{1}{M}\sum_{k=1}^{M}F_{k}(\overline{\mathbf{W}})$ where $\overline{\mathbf{W}}=\mathbf{W}_{0}+\overline{\mathbf{B}}\leavevmode\nobreak\ \overline{\mathbf{A}}$. Throughout the paper, we denote the truncation of a matrix with the :::: symbol for each row and column at the subscript. For instance, if there is no truncation at the row but truncation from $x<l$ at the column for the matrix $\mathbf{W}\in\mathbb{R}^{d\times l}$, we keep all the columns until $x$ and omit the last $l-x$ columns and denote the resulting matrix it as $\mathbf{W}_{:,:x}$.
3.2 Naïve Case: Homogeneous LoRA
A straightforward way to perform federated fine-tuning with LoRA is to train the LoRA modules $\mathbf{B},\leavevmode\nobreak\ \mathbf{A}$ with homogeneous rank $r$ across all clients with standard FL McMahan et al. (2017). Specifically, first the clients have the pre-trained ODFM weights $\mathbf{W}_{0}$ stored in their devices prior to training for the forward pass when training the LoRA modules. Then, the server sends the global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ to the set of $m$ selected clients $\mathcal{S}^{(t)}$ per communication round $t$. Each selected client $k\in\mathcal{S}^{(t)}$ trains the LoRA modules on their local data for a few local iterations (usually with mini-batch SGD) and send the updated modules $\mathbf{B}_{k}^{(t)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t)}$ back to the server. The server then finally updates the global LoRA modules accordingly to $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{B}_{k}^{(t)}/m,\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\mathbf{A}_{k}^{(t)}/m$ and send back to the next set of selected clients for the next communication round. This training process is nearly identical to the standard FL algorithm McMahan et al. (2017) except that the pretrained weights $\mathbf{W}_{0}$ are freezed and locally stored in the clients’ devices and only the LoRA moduels are trained and communicated.
Instead of such homogeneous rank deployment, however, for federated fine-tuning it is not only possible but more practical to consider heterogeneous rank deployment where each client trains different rank LoRA modules. Such setting is motivated and often required from the system constraints of the clients Wang et al. (2021) where most of the clients are only capable of having smaller ranks while a few can handle larger ranks. However, with heterogeneous ranks across for clients, it is non-trivial how to aggregate the heterogenenous LoRA modules at the server and then re-distribute them back to the next set of selected clients. Pushing the limits beyond homogeneous LoRA, we propose HetLoRA in the next section.
Figure 1: Example of (a): the pretrained weights $\mathbf{W}_{0}$ that are stored in the clients’ devices and (b): the heterogeneous ranks assigned to different clients with $r_{\text{min}}=r_{1}<r_{2}<r_{3}=r_{\text{max}}$ for heterogeneous LoRA.
Figure 2: Overview of the zero-padding, sparsity-weighted aggregation, and truncation method for HetLoRA; (a): Zero-pad LoRA modules with smaller ranks to $r_{\text{max}}$ (clients with rank $r_{\text{max}}$ does not need padding) and calculate their sparsity by calculating the Frobenius norm of the reconstructed model $\Delta\mathbf{W}_{k}^{(t)}=\mathbf{B}_{k}^{(t)}\mathbf{A}_{k}^{(t)}$; (b): After padding, aggregate all of the clients’ LoRA modules with the weights $\|\mathbf{S}_{k}^{(t)}\|/\mathbf{Z}^{(t)}$ calculated by $\Delta\mathbf{W}_{k}^{(t)}$ to get the global LoRA modules; (c): Truncate the global LoRA modules for the specific rank of the next selected client (example for client with rank $r_{2}$).
3.3 Proposed Method: Heterogeneous LoRA
Overview. Recall in heterogeneous LoRA, different clients have different ranks initially assigned to them prior to training to match their computation resources. Our proposed heterogeneous LoRA method, namely HetLoRA, does not assume any specific method to assign the ranks to the clients since we believe that the ranks are not to be decided by the server but rather by the clients and their available resources. We formulate that each client $k$ has a rank $r_{k}$ that is initially assigned to the client where the ranks range from a minimum and maximum rank, i.e., $r_{k}\in[r_{\text{min}},\leavevmode\nobreak\ r_{\text{max}}],\leavevmode\nobreak\ \forall k$ (see Fig. 1). In our proposed HetLoRA there are three major steps for federated fine-tuning of heterogeneous rank LoRA modules: 1) Distribution via Truncation, 2) Local Training with Rank Self-Pruning, and 3) Sparsity-Weighted Aggregation of the LoRA modules. We go into the details of each step in the subsequent paragraphs in this section. An overview of HetLoRA is illustrated in Fig. 2.
1) Distribution via Truncation. First, for each communication round $t$ the server has its initial global LoRA modules $\overline{\mathbf{B}}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}$ with a global rank $r^{(t)}$. The value of the global rank $r^{(t)}$ depends on how we aggregate the heterogeneous rank LoRA modules which we will elaborate on in step 3). The server has to distribute these global LoRA modules to a subset of selected set of clients $\mathcal{S}^{(t)}$ with heterogeneous ranks $r_{k}^{(t)},k\in\mathcal{S}^{(t)}$ for local training [^1] There is a superscript $t$ for the ranks $r_{k}^{(t)}$ across clients which indicates that in HetLoRA these heterogeneous ranks can be changed over the communication rounds via self-pruning explained in step 2.. With the given global LoRA modules, we consider a simple and intuitive method of truncation where the server sends $\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t)}_{:r_{k},:}$ to each client $k$ with rank $r_{k}^{(t)}$ for local training where we omitted the superscript for $r_{k}$ for simplicity.
2) Local Training with Rank Self-Pruning. Initializing their local LoRA modules with the ones received from the server as $\mathbf{B}_{k}^{(t,0)}=\overline{\mathbf{B}}^{(t)}_{:,:r_{k}},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,0)}=\overline{\mathbf{A}}^{(t)}_{:r_{k},:}$, each client $k\in\mathcal{S}^{(t)}$ can perform $\tau$ local iterations of mini-batch SGD on their local data to minimize the local objective $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$ to send back the updated LoRA modules $\mathbf{B}_{k}^{(t,\tau)}\in\mathbb{R}^{d\times r_{k}^{(t)}}$ and $\mathbf{A}_{k}^{(t,\tau)}\in\mathbb{R}^{r_{k}^{(t)}\times l}$. This is the same process as the standard local training step in vanilla FedAvg McMahan et al. (2017). However, we improve this vanilla local training step by adding a rank self-pruning mechanism where clients self-prune their respective ranks depending on the magnitude of the model parameters.
Specifically, we add a regularization term to the original local objective to get $\min_{\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}}\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})+\lambda\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ where $\gamma<1$ is a decay-factor that determines how aggressively we want to prune the ranks to a smaller value. The regularization term aims to minimize the norm of the last few ranks, which will become smaller if the first loss term $\frac{1}{|\mathcal{B}_{k}|}\sum_{\xi\in\mathcal{B}_{k}}\ell((\mathbf{B}_{k},\leavevmode\nobreak\ \mathbf{A}_{k}),\xi|\mathbf{W}_{0})$ is not very large. After training with the new local objective we compare the norm of the updated LoRA modules’ last few layers $\left\|\mathbf{B}_{k,:,r_{k}\gamma:r_{k}}\right\|\left\|\mathbf{A}_{k,r_{k}\gamma:r_{k},:}\right\|$ with the ones from the initially received LoRA modules. If the former is smaller we prune the last few layers (which pruning intensity is determined by $\gamma$) and send back the LoRA modules with a smaller rank. This means that for the LoRA modules which incurs a small local loss, i.e., well-trained on the clients’ local data, the LoRA modules are more likely to be pruned to a smaller rank. Such pruning allows HetLoRA to reduce the noise in the LoRA modules introduced by clients having a larger rank than the actual rank that their data complexity requires and also reduce the model complexity of the LoRA modules to improve generalization and prevent overfitting (see Table 4). Once the rank is pruned for a client, the client saves the updated rank and uses it as the starting rank if selected for future communication rounds. The client then sends back their updated and possibly rank-pruned LoRA modules to the server for the modules to be processed in the next step, server-side aggregation.
3) Sparsity-Weighted Aggregation. Finally, the last step of HetLoRA is aggregating the received heterogeneous LoRA modules $\mathbf{B}_{k}^{(t,\tau)},\leavevmode\nobreak\ \mathbf{A}_{k}^{(t,\tau)},k\in\mathcal{S}^{(t)}$. A straightforward way to aggregate the hetergeneous LoRA modules is using zero-padding to all the received LoRA modules with $r_{i}^{(t)}<\max\{r_{k}^{(t)}|k\in\mathcal{S}^{(t)}\}$ and then perform simple averaging over the modules. However, such naive aggregation can lead to biasing the model towards higher rank clients even when these clients may not hold valuable training information, i.e., have low data complexity, giving noisy updates.
In an ideal scenario where we can deploy any rank to any client, deploying higher ranks to the clients with higher data complexity or larger local datasets can retrieve more informative and less sparse updates from the clients. Conversely if we assign higher ranks to the clients whose data complexity is low, the actual rank of the full model from the reconstructed LoRA modules can be smaller than the assigned rank. Thus the higher rank client’s update may be unnecessarily over-emphasized in the naive zero padding method.
Based on this insight we propose a sparsity-weighted aggregation scheme where the server reconstructs these LoRA modules to the full model as $\Delta\mathbf{W}_{k}^{(t)}=\mathbf{B}_{k}^{(t)}\mathbf{A}_{k}^{(t)}$ and gets the norm of the singular value vectors from the full models denoted as $\mathbf{S}_{k}^{(t)}$ by calculating $\|\Delta\mathbf{W}_{k}^{(t)}\|_{F}$. Note that the costly process of performing SVD for each of the full model $\Delta\mathbf{W}_{k}^{(t)}$ can be avoided by simply calculating the Frobenius norm of $\Delta\mathbf{W}_{k}^{(t)}$ (see Lemma 1.2 in Guruswami and Kannan (2012)). The server then weighs the LoRA modules accordingly by the norm of the singular value vectors, updating the global LoRA modules as $\overline{\mathbf{B}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|\mathbf{B}_{k}^{(t)}/\mathbf{Z}^{(t)},\leavevmode\nobreak\ \overline{\mathbf{A}}^{(t+1)}=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|\mathbf{A}_{k}^{(t)}/\mathbf{Z}^{(t)}$ where $\mathbf{Z}^{(t)}\vcentcolon=\sum_{k\in\mathcal{S}^{(t)}}\|\mathbf{S}_{k}^{(t)}\|$ is the normalizing factor. This way, we can de-emphasize the larger rank assigned-clients that have rather less informative updates, and more emphasize the smaller rank assigned-clients that have more informative ones.
3.4 Why not Simply Reconstruct First,
then Redistribute the LoRA modules?
One might ask why not simply reconstruct each of the LoRA modules to the full matrix and aggregate them. Here we show that reconstructing the LoRA modules and aggregating them to get the full model results in a different full model compared to when we aggregate the LoRA modules first and then reconstruct the final model. In Section 4 we also empirically show that reconstructing the LoRA modules to the full model and redistributing them after truncated SVD to the corresponding rank of the clients results in an underwhelming performance compared to HetLoRA.
[…]
One can observe that the difference between the two models are the cross-terms between the left and right module of different client 1 and 2, i.e., $b_{1,i0}a_{2,0j}+b_{2,i0}a_{1,0j}$ for the $i^{th}$ row and $j^{th}$ column. In other words, when we reconstruct the LoRA modules first and then aggregate them to get the full model, each term in the full model are cross-products between the left and right module of each client and not the cross-products between clients. Thus, reconstructing the LoRA modules loses information on the cross-relation across clients, only retaining the knowledge on the cross-relation between the LoRA modules $\mathbf{B}$ and $\mathbf{A}$. Such observation is also corroborated by the reconstruction first’s underwhelming performance in Table 3.
Table 3: Final RougeL score for Reddit text summarization and perplexity for multi-session chat for different federated fine-tuning methods. The blue text indicates the ratio of trained number of parameters compared to the full-fine tuning case. HetLoRA outperforms both HomLoRA and Recon+SVD method, but slightly underperforms the full-fine tuning case. However, compared to full fine-tuning the number of trained parameter is significantly smaller.
Reddit (RougeL) Multi-Session Chat (Perplexity)
PaLM-XXS PALM-XS PaLM-XXS PALM-XS
Full $\mathbf{94.56}(\pm 0.01)$ $\mathbf{94.87}(\pm 0.04)$ $\mathbf{32.70}(\pm 0.17)$ $\mathbf{23.40}(\pm 0.36)$
[…]
$\textsc{HetLoRA}\leavevmode\nobreak\ \gamma=0.99$ $\mathbf{94.23}(\pm 0.03),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\times 0.00}3$ $\mathbf{94.41}(\pm 0.05)$ $\mathbf{53.93}(\pm 1.57),{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\times 0.002}$ $\mathbf{38.76}(\pm 0.52)$
4 Experiments
In this section, we present results for HetLoRA and its baselines in terms of the performance on training speed, computation/communication efficiency, and final achieved performance. First, we show the performance of homogeneous LoRA to show how LoRA in general performs for low and high rank values. Second, we demonstrate HetLoRA’s performance for different $r_{\text{min}}$ and $r_{\text{max}}$ values comparing them with full-fine tuning, homogeneous LoRA, and the reconstruction-first method elaborated in Section 3.4. We also conduct an ablation study on HetLoRA with varying decay factor $\gamma$ for the rank self-pruning step.The rank distribution across clients for HetLoRA, unless mentioned otherwise, is set to a truncated power-law distribution with $\alpha=0.1$ in the range between $[r_{\text{min}},r_{\text{max}}]$ (inclusively), where the small $\alpha$ value makes the distribution skewed towards smaller ranks. All experiments were ran with 3 different random seeds and their average is shown along with the standard deviation.
Model. We use the transformer-based model PaLM2 Google (2023) of size XXS and XS for our experiments which we consider applicable to fit in to the category of ODFMs given their smaller sizes (340M-1B number of parameters) compared to standard FMs. The LoRA modules are applied to only the self-attention layers as proposed in the original LoRA paper Hu et al. (2021), and their relative number of parameters compared to the original model are shown in Table 2.
Tasks. The tasks we consider are the chat dialogue from the multi-session chat (MSC) dataset Xu et al. (2021) and the text summarization task from the Reddit dataset Völske et al. (2017). The MSC data is a collection of human-human interactions comprising numerous extended chat sessions, and we use perplexity Zhang et al. (2018) as the metric which has been used to show the quality of chat responses from generative models from previous literature Sedoc et al. (2019). We sample 100 users uniformly at random and partition their data for training and evaluation by each previous_dialogs and dialog. The Reddit text summarization data consists of real users’ reddit posts and their summarization, and we use RougeL Lin (2004) as the metric. We use 298 users from Reddit that have at least 100 data samples as the training clients and use another 100 users with at least 100 data samples for evaluation.
[…]
(a) PaLM-XXS
(b) PaLM-XS
Figure 3: Performance of homogeneous LoRA for different rank $r$. Higher ranks achieve better performance with fewer communication rounds than the lower ranks, but they overfit quickly. Conversely, the lowest rank $r=1$ achieves low perplexity slower than higher ranks, but without overfitting.
4.1 Experiment Results
Homogeneous LoRA and the Effect of Ranks $r$. First, we evaluate the performance of federated fine-tuning of the LoRA modules with homogeneous LoRA deployment across clients in Fig. 3 for different ranks $r\in[1,5,20,50]$. We observe that a higher rank $r$ for homogeneous LoRA achieves better perplexity floor with fewer communication rounds than the lower ranks but quickly overfits resulting in worse performance compared to the lower ranks after more communication rounds. On the other hand, while the lower rank cases need more communication rounds to achieve good performance, it does not have the problem of overfitting as the higher ranks. Hence for homogeneous LoRA, there is a trade-off to consider between low and high ranks, in terms of faster performance achievement and overfitting. Note that these observations are consistent with previous literature in the centralized setting where a higher rank does not necessarily yields the best performance Hu et al. (2021); Zhang et al. (2023b). Next, we show that HetLoRA achieves good performance quickly without this overfitting issue, showing better performance than the homogeneous LoRA case.
(a) $r_{\text{max}}=10$
(b) $r_{\text{max}}=50$
Figure 4: Performance of HetLoRA without rank pruning or and with simple average aggregation. Similar to homogeneous LoRA, larger $r_{\text{min}}$ leads to overfitting for heterogeneous LoRA, but it is not as severe as homogeneous LoRA even for larger maximum rank $r_{\text{max}}=50$ showing that the smaller rank LoRA modules act as a regularizer for HetLoRA.
Naïve Heterogeneous LoRA and the Effect of $r_{\text{min}}$ and $r_{\text{max}}$. First, we show the performance of naïve heterogeneous LoRA without self rank-pruning and with only average aggregation instead of the sparsity-weighted aggregation in Fig. 4. We can see similar observations to those from homogeneous LoRA where a smaller minimum rank $r_{\text{min}}=1$ leads to slower training but better performance while a larger maximum rank leads to faster training but worse performance. However, compared to homogeneous LoRA the overfitting does not get as severe for heterogeneous LoRA even with much larger ranks such as $r_{\text{max}}=50$. We can imply from this result that the smaller rank LoRA modules act as a regularizer in heterogeneous LoRA. Next, we show that by adding the self rank-pruning and sparsity-weighted aggregation, even with $r_{\text{min}}=5$ we are able to prevent overfitting issues and achieve better training speed and final performance than other baselines.
Table 4: Ablation study on the effect of the decaying factor $\gamma$ for HetLoRA’s self-rank pruning in the local training step. While aggressive pruning can be harmful to HetLoRA’s performance, pruning ($\gamma=0.99$) can outperform the case when there is no pruning at all ($\gamma=1$) by reducing the noise introduced by large rank clients with low data complexity.
Reddit (RougeL) Multi-Session Chat (Perplexity)
PaLM-XXS PALM-XS PaLM-XXS PALM-XS
$\gamma=1$ $92.17(\pm 0.08)$ $91.95(\pm 0.03)$ $55.07(\pm 0.81)$ $40.92(\pm 0.58)$
$\gamma=0.99$ $\mathbf{94.23}(\pm 0.03)$ $\mathbf{94.41}(\pm 0.05)$ $\mathbf{53.93}(\pm 1.57)$ $\mathbf{38.76}(\pm 0.52)$
$\gamma=0.95$ $89.62(\pm 1.33)$ $83.19(\pm 1.70)$ $71.10(\pm 1.39)$ $46.39(\pm 0.87)$
$\gamma=0.85$ $60.31(\pm 3.04)$ $53.28(\pm 2.47)$ $120.72(\pm 10.93)$ $59.67(\pm 1.98)$
Heterogeneous LoRA compared to Baselines. Finally, we compare our proposed HetLoRA with other baselines in Table 3 and Fig. 5. We see that HetLoRA with $r_{\text{min}}=5$ and $r_{\text{max}}=50$ achieves faster training as well as better performance than homogeneous LoRA cases with both edge cases of the ranks $r\in\{5,50\}$ and reconostruction+SVD which was explained in Section 3.4. This implies that HetLoRA is not only practical in the sense that clients are allowed to have their own rank values, it can also outperform the limited case of homogeneous LoRA where all clients have $r=r_{\text{min}}$ or the impractical case where all clients have $r=r_{\text{max}}$. We also observe that HetLoRA achieves slightly lower performance than full fine-tuning. However, as shown in the blue text in Table 3 that shows the number of trained parameters compared to the full-fine tuning case, full fine-tuning requires to train a much larger number of parameters compared to HetLoRA, making it infeasible to train with ODFMs in practice. We also show in Fig. 6 that to achieve the targeted performance for both Reddit and MSC task, HetLoRA requires significantly less number of parameters to be trained and communicated compared to full-fine tuning. Although for Reddit, HomLoRA has a slightly less number of parameters to be trained, the final achieved RougeL is outperformed by HetLoRA as shown in Table 3.
(a) Reddit
(b) MSC
Figure 5: Comparison of the performance across homogeneous LoRA, heterogeneous LoRA, and full-fine tuning. Heterogeneous LoRA achieves better performance than homogeneous LoRA with fewer number of communication rounds.
Figure 6: Ratio of communicated number of parameters for different PEFT methods to full fine-tuning to achieve the target value for the metric where it is RougeL 80 for Reddit text summarization task and perplexity 150 for the multi-session chat response task. The ‘X’ means that the target metric is not achieved even after convergence.
Effect of the Decaying Factor $\gamma$. Lastly, we conduct an ablation study on the effect of the decaying factor $\gamma$ of HetLoRA’s local training step with self-rank pruning in Table 4. We observed that aggressive pruning hurts the performance where $\gamma=0.85$ shows the worse performance across the varying $\gamma$ values. On the other hand, no pruning at all ($\gamma=1$) underperforms the case when there is pruning ($\gamma=0.99$), showing that reducing the noise introduced by large rank clients which data complexity is actually not that high indeed improves the performance.
5 Discussions and Concluding Remarks
In our work, we investigated federated fine-tuning for ODFMs that cater to device system and data heterogeneity with our proposed HetLoRA. We show that HetLoRA is not only practical but also achieves better training speed, communication/computation efficiency, and final performance compared to homogeneous LoRA. We also show interesting findings consistent with previous literature Hu et al. (2021); Zhang et al. (2023b) that increasing ranks does not always help for homogeneous LoRA. Our findings in this work opens up several questions worth investigating. For instance, if the settings allow us to assign specific ranks to clients what will be the effective way to assign the ranks across clients for better convergence and performance? Another important next step of our work includes pursuing the theoretical convergence and generalization of heterogeneous LoRA.